# Solving quantum-inspired dynamics on quantum and classical annealers

### v1

@@ Acknowledgements
The authors acknowledge the Jülich Supercomputing Centre for providing computing time on the D-Wave Advantage™ System JUPSI through the Jülich UNified Infrastructure for Quantum computing (JUNIQ). P.H. and B.G. acknowledges funding from the National Science Centre (NCN), Poland, under Projects: Sonata Bis 10, No. 2020/38/E/ST3/00269. Z.M. acknowledges funding from the Ministry of Economic Affairs, Labour and Tourism Baden-Württemberg in the frame of the Competence Center Quantum Computing Baden-Württemberg (project “KQCBW24”).
@@
Philipp Hanussek1,2, Jakub Pawłowski 3,4, Zakaria Mzaouali5, Bartłomiej Gardas1
@@
Figure S1: Plots of system-resolved $\text{TTS}_{99}$ as a function of system size, demonstrating clear exponential scaling $\propto\exp(N/\beta)$, with $\beta$ exponent indicated in the legend. Missing data points indicate zero success probability, which is the case e.g. for quantum solver in systems $H_{3}$ and $H_{8}$. In the $N\gtrsim 100$ regime the advantage of GPU solvers becomes clearly visible. Nevertheless, the new Advantage2 1.6 QPU demonstrates the developing potential of quantum annealaers.

### v2

@@ Acknowledgements
The authors acknowledge the Jülich Supercomputing Centre for providing computing time on the D-Wave Advantage™ System JUPSI through the Jülich UNified Infrastructure for Quantum computing (JUNIQ).
P.H. and B.G. acknowledges funding from the National Science Centre (NCN), Poland, under Projects: Sonata Bis 10, No. 2020/38/E/ST3/00269. Z.M. acknowledges funding from the Ministry of Economic Affairs, Labour and Tourism Baden-Württemberg in the frame of the Competence Center Quantum Computing Baden-Württemberg (project “KQCBW25”).
@@
Philipp Hanussek1,2, Jakub Pawłowski 3,4, Zakaria Mzaouali5,6, Bartłomiej Gardas1
@@
6
Jülich Supercomputing Centre, Institute for Advanced Simulation, Forschungszentrum Jülich, Wilhelm-Johnen-Straße, Jülich, 52428, Germany.
@@
Figure S1: Plots of system-resolved $\text{TTS}_{99}$ as a function of system size, demonstrating clear exponential scaling $\propto\exp(N/\beta)$, with $\beta$ exponent indicated in the legend. Missing data points indicate zero success probability, which is the case e.g. for quantum solver in systems $H_{3}$ and $H_{8}$. In the $N\gtrsim 100$ regime the advantage of GPU solvers becomes clearly visible. Nevertheless, the new Advantage2 1.6 QPU demonstrates the developing potential of quantum annealers.